### v1

@@ 4 Experiments
In this section, we aim to answer the following questions:
  1. Q1.
    Quantitative Comparison: How well does TSAP perform on TSAD against established baselines?
  2. Q2.
    Qualitative Analysis: How well does TSAP tune the augmentation hyperparameters?
  3. Q3.
    Ablation Studies: Which design decisions in TSAP drive its effectiveness?
4.1 Setup
Datasets
We evaluate TSAP on six distinct TSAD tasks. Four of these are conducted in a controlled environment, while the remaining two are natural. In the controlled setting, the anomaly types are manually injected in $\mathcal{D}_{\mathrm{test}}$ based on the types discussed in Sec. 3.1. The controlled environment allows for a thorough assessment of TSAP’s ability to perform continuous hyperparameter tuning. For the natural environment, the anomaly types in $\mathcal{D}_{\mathrm{test}}$ are a priori unknown and it is the goal of TSAP to find the anomaly type that yields best alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ (part of $\mathcal{D}_{\mathrm{test}}$), expressed by $\mathcal{L}_{\mathrm{val}}$. The details of the datasets and anomaly types are provided in Appx. 0.A.1 and 0.A.2.
For controlled tasks, we use the 2017 PhysioNet Challenge dataset [9], which comprises real-world ECG recordings. Table 1 shows a selection of four different controlled TSAD tasks, constructed by manually injecting different anomaly profiles into the PhysioNet dataset. For PhysioNet A and B, given the anomaly type (Platform), the task is to infer or tune, respectively, the hyperparameter(s) level only and both level and length, while anomaly location is random (hence not tuned). For PhysioNet C and D, the respective tuning tasks are the same but for a different anomaly type (Trend). We choose Platform and Trend anomalies to represent, respectively, range-bound and range-exceeding anomalies. Additional variations of PhysioNet with other injected anomaly types, such as Mean shift and Extremum, are discussed in Appx. 0.A.2.
The natural TSAD tasks are derived from the CMU Motion Capture (MoCap) dataset [^2] http://mocap.cs.cmu.edu/. We consider the walking signal as normal data, and the signals of jumping and running as anomalies. To generate normal signals, we stitch the walking signals by identifying the start and end points of each gait phase and add random noise; whereas to generate anomalous ones, we stitch walking or running signals at a random location in the normal signal. This yields two distinct TSAD tasks as shown in Table 1. Different from PhysioNet A–D where we only tune the continuous hyperparameter(s) for the given (discrete) anomaly type, here for MoCap A and B, we aim to tune both the unknown anomaly type that corresponds to Jump and Run behavior, respectively, as well as the (continuous) hyperparameter level while location and length take random values.
Table 1: Anomaly profile of different TSAD tasks.
Profile PhysioNet A PhysioNet B PhysioNet C PhysioNet D MoCap A MoCap B
Type Platform Platform Trend Trend Jump Run
Level Fixed Fixed Fixed Fixed Fixed Fixed
Location Random Random Random Random Random Random
Length Random Fixed Random Fixed Random Random
Baselines
We compare TSAP with a selection of established baselines, including both traditional and deep learning-based methods that have demonstrated efficacy in TSAD [29]. Among the traditional methods, we utilize three with different modeling approaches; namely, One-Class Support Vector Machines (OC-SVM) [30]; Local Outlier Factor (LOF) [7]; and Isolation Forest (IF) [21]. On the deep learning side, we benchmark against the Encoder-Decoder LSTM (EncDec-LSTM) [24]; the Spectral Residual Convolutional Neural Network (SR-CNN) [28]; and the Unsupervised Anomaly Detection (USAD) for TSAD [2]. Lastly, we include a state-of-the-art competing method which learns augmentations in the embedding space, called Neural Transformation Learning for (TS)AD (NeuTraL-AD) [26]. This diverse set of baselines allows for a comprehensive analysis across different approaches within the TSAD domain.
Model Configurations
The Encoderϕ in $f_{\mathrm{aug}}$ and Encoderθ in $f_{\mathrm{det}}$ are constructed using 1D CNN blocks [4] (transposed 1D CNN for Decoderϕ) for efficient temporal feature extraction. We carefully choose the number of epochs $T$ to allow sufficient time for the convergence of $\mathbf{a}$, with empirical evidence suggesting that $T=100$ typically suffices. For the number of inner-loops $L$, we opt for $L=5$, aligned with [34], such that $f_{\mathrm{det}}$ has adequate time to learn effective discriminative embeddings for $\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{trn}}$. A detailed overview of the configuration for TSAP and all baselines is provided in Appx. 0.A.3.
Evaluation Metrics
Our method calculates anomaly scores on an entire sequence level $\mathbf{x}$, similar to [26]. This is a different set-up compared to novelty detection in time series which typically operates on a point level. Detection on a sequence level can be especially important to spot long-term patterns (e.g. Trend anomalies). As such, we use the Area Under the Receiver Operating Characteristic Curve (AUROC) and the $F_{1}$ score as key performance metrics to quantify detection capability of anomalous sequences. All results are reported on the unseen $\mathcal{D}_{\mathrm{test}}$. In addition, we determine the optimal $F_{1}$ score, by enumerating all possible thresholds, given by the anomaly scores for a given segment $\mathbf{x}$. We then compute the corresponding precision and recall for each threshold and select those that yield the highest $F_{1}$. As such, AUROC provides a balanced view whereas $F_{1}$ shows the optimal scenario. Both metrics range from 0 to 1, with higher values indicating superior performance.
4.2 Quantitative Results
Table 2: Detection peformance of baselines w.r.t. $F_{1}$ and AUROC on test data across six TSAD tasks. Four tasks are performed in a controlled setting (anomaly type is manually injected and corresponds to anomaly types described in Sec. 3.1), based on the PhysioNet ECG data. Remaining two TSAD tasks exhibit natural anomalies (real-world anomaly type which is unknown), based on the MoCap data. TSAP outperforms most baselines and has the best (lowest) average rank w.r.t. both $F_{1}$ and AUROC, with low standard deviation (in parentheses).
@@ 4.2 Quantitative Results
OC-SVM 0.182 0.468 0.182 0.472 0.373 0.803 0.393 0.806 1.000 1.000 0.546 0.806 5.7 (2.6) 5.3 (2.3)
LOF 0.999 1.000 0.999 1.000 0.354 0.738 0.358 0.725 0.196 0.506 0.221 0.577 5.5 (2.7) 5.2 (3.3)
IF 0.255 0.587 0.232 0.576 0.183 0.402 0.182 0.356 0.864 0.965 0.342 0.758 6.2 (1.6) 6.3 (1.4)
EncDec-LSTM 0.190 0.508 0.190 0.508 0.415 0.812 0.442 0.819 0.980 0.999 0.909 0.996 4.8 (1.0) 5.00 (0.9)
SR-CNN 0.965 0.990 0.964 0.998 0.983 0.999 0.991 0.999 0.302 0.700 0.214 0.512 3.8 (3.0) 4.0 (2.8)
USAD 0.183 0.425 0.184 0.428 0.430 0.822 0.409 0.828 1.000 1.000 1.000 1.000 4.2 (2.7) 4.3 (3.1)
NeuTraL-AD 0.211 0.732 0.263 0.679 0.561 0.868 0.526 0.862 1.000 1.000 1.000 1.000 2.8 (1.6) 2.7 (1.4)
TSAP (ours) 1.000 1.000 1.000 1.000 0.973 0.999 0.991 0.998 0.889 0.969 1.000 1.000 1.8 (1.6) 1.8 (1.6)
@@
Table 2 provides the detection results for all six TSAD tasks described in Sec. 4.1. TSAP ranks the best overall in terms of both $F_{1}$ and AUROC. This shows that detector $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$, trained through the alternating mechanism of TSAP, is able to generalize to unseen and unlabeled anomalies in $\mathcal{D}_{\mathrm{test}}$.
While some competing methods perform strongly on a subset of tasks, they lack consistency across all TSAD tasks. Among the traditional baselines, which perform subpar as compared to deep TSAD approaches, LOF performs very well in PhysioNet A and B, relatively poorly on C and D, but exhibits near-random performance on MoCap. This discrepancy can be attributed to the nature of the anomalies in the PhysioNet data, which are manually injected and often characterized by pronounced and abrupt changes as shown in Fig. 2. These abrupt changes lead to large differences in local density, which makes LOF thrive. In contrast, the MoCap dataset contains anomalies that emerge more subtly, as one gradually transitions from walking to running or jumping, reflecting more natural variations in the data which, in turn, lead to more subtle differences in local density. Similarly, SR-CNN performs strongly on the PhysioNet TSAD tasks, yet its effectiveness also diminishes drastically on MoCap tasks. SR-CNN exploits the abrupt changes present in the PhysioNet data but fails to detect the more subtle anomalies in MoCap. This suggests that, while SR-CNN and LOF work well for certain types of anomalies, their utility may be limited in scenarios where anomalies are more subtle.
On the contrary, reconstruction-based baselines such as EncDec-LSTM and USAD face challenges on PhysioNet due to its high variability among inliers, which complicates the task of accurate data reconstruction. Yet, these methods excel with the MoCap dataset, where its consistent near-periodic pattern (Fig. 2, bottom) lends itself to more reliable reconstruction, thereby enhancing anomaly detection performance. Similarly, NeuTraL-AD, a state-of-the-art augmentation-based baseline, demonstrates proficiency in properly augmenting the inlier data within MoCap. However, its performance on the PhysioNet variations reveals some weaknesses. NeuTraL-AD struggles with the high variability in the PhysioNet inliers, inherent to real-world ECG signals. This suggests that the augmentation functions considered in NeuTraL-AD lack robustness when inliers are inherently noisy.
We remark that only TSAP provides robust and consistent performance across all TSAD tasks, showcasing the effectiveness and generalizability of our proposed method.
4.3 Qualitative Results
@@
Figure 3: (best in color) Tuning continuous augmentation hyperparameter(s) with TSAP. Top: Given Platform anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val. loss (center), and leading to high detection performance (right). Bottom: Multiple continuous hyperparameters, here both level and length, are accurately tuned to near true values (left), as guided by minimizing the val. loss (center), achieving high AUROC (right). Notice that the diverged results in both cases associate with high (or hard-to-optimize) val. loss, which help us effectively reject low performance models.
A key contributor to TSAP’s consistent performance is the validation loss through which TSAP automatically learns the augmentation hyperparameters $\mathbf{a}$. Once $\mathbf{a}$ is determined, the task essentially reduces to a supervised learning problem. Next, we show that TSAP not only effectively tunes the continuous augmentation hyperparameters $\mathbf{a}$, but also that its validation loss guides the accurate selection of the discrete hyperparameter (i.e. anomaly type).
Controlled Environment
Consider PhysioNet A, where we aim to tune the continuous hyperparameter $a$, i.e. level, of the Platform anomalies present in $\mathcal{D}_{\mathrm{test}}$. That is, the level in $\mathcal{D}_{\mathrm{test}}$ is fixed and tuning aims to estimate its value using TSAP while the other hyperparameters (location, length) are randomized. Fig. 3 (top) shows TSAP’s estimation process for different initializations of $a$. We observe that the initialization for $a\in\{-0.4,0.6\}$ leads to the true $a$$=$$0.2$ (left). Simultaneously, the validation loss drops substantially once TSAP has arrived at the true $a$ (middle). This is also reflected in the performance of $f_{\mathrm{det}}$ on part of the unlabeled test data $\mathcal{D}_{\mathrm{val}}$ which soars upon estimation of the true $a$ (right). Conversely, the initialization for $a\in\{0.8,-0.8\}$ leads to a high validation loss, indicating poor alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$. Indeed, the performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ now suffers from poor alignment.
For PhysioNet B, we estimate both level and length while location is randomized. Fig. 3 (bottom) demonstrates TSAP’s ability to accurately estimate the level and length. While this is a by-product of our method, there are several real-world use-cases that can directly benefit from accurately learning the anomaly profile, from industrial equipment monitoring, to network security, and healthcare monitoring. This demonstrates TSAP’s versatility given its capability of estimating continuous hyperparameters. See Appx. 0.A.4 for additional qualitative results on other anomaly types such as Trend, Mean shift, and Spike.
Further, Fig. 4 (top) showcases TSAP’s ability to perform discrete hyperparameter selection. Here, TSAP has been initialized and trained with three different anomaly types (Mean shift, Platform, Trend) on PhysioNet C (true anomaly type is Trend). The validation loss clearly showcases a misalignment between the Platform and Trend types (left), which is also reflected in the AUROC of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ (middle). Note how the Mean shift anomaly type has a low $\mathcal{L}_{\mathrm{val}}$ towards the end of the training epochs, which is also reflected in the AUROC on $\mathcal{D}_{\mathrm{val}}$. Indeed, comparing TSAP tuned for Mean shift anomalies and for Trend anomalies show strong resemblances with the true underlying anomaly type (right). This demonstrates that the true underlying anomaly type is not necessarily the only type that yields high alignment, and in turn, a high-performing detector.
Natural Environment
In MoCap datasets, the anomaly types are a priori unknown. As such, we initialize TSAP with different augmentation types (Frequency and Platform) to perform discrete hyperparameter selection. Fig. 4 (bottom) highlights its effectiveness as the validation loss clearly prefers one type over the other. Indeed, the natural anomalies defined by jumping signals in MoCap A have close resemblance to platform anomalies. Discrete hyperparameter optimization for MoCap B is similar, as given in Appx. 0.A.4.
@@
Figure 4: (best in color) Tuning the discrete hyperparameter (anomaly type) with TSAP. Top: Given true Trend anomaly (black), val. loss favors both Trend (green) and Mean shift (blue) type, both with high AUROC (middle) and high resemblance (right), and effectively rejects type Platform (orange) with poor performance. Bottom: For Jump anomalies in MoCap A with unknown type (black), val. loss favors type Platform (purple) that leads to high AUROC (middle) and mimics well the true anomaly (right), and effectively rejects type Frequency (red) with poor performance.
4.4 Ablation Studies
We showcase two ablation studies on the controlled PhysioNet data to support various design strategies of TSAP. Additional ablation studies on embedding normalization and second-order optimization are provided in Appx. 0.A.4.
Validation Loss
Figure 5: (best in color) Overview of ablation studies. Top: TSAP’s $\mathcal{L}_{\mathrm{val}}$ is replaced with a point-wise val. loss leading to an erroneous estimation of $a$ (left) and poor performance (right). Bottom: TSAP’s self-tuning module is disabled and $a$ is instead randomized. Val. loss indicates poor alignment, reflected in poor performance.
Randomization vs. Tuning
In Fig. 5 (bottom), the self-tuning module is disabled for PhysioNet C, where level is instead randomized (along with location, and length). We observe substantially higher $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment. In turn, $f_{\mathrm{det}}$ struggles to detect the unlabeled anomalies in $\mathcal{D}_{\mathrm{val}}$. This clearly demonstrates the utility of TSAP’s systematic hyperparameter (self-) tuning over random choice.
5 Related Work
Self-supervised Learning (SSL) for Anomaly Detection:
SSL has emerged as a significant approach in machine learning. Foundation models [6] like large language models [25, 27], whose training heavily relies on SSL, have shaken up the world with outstanding performance. SSL has also transformed representation learning and significantly boosted several tasks in NLP, computer vision, recommender systems and medicine [3, 19]. SSL is especially attractive for unsupervised problems like anomaly detection (AD) [16], because of its nature to create proxy tasks and loss functions without any labels. There are various types of SSL-based methods on AD [22], but the core idea is to use data augmentation functions [8, 14, 20] on inliers to create pseudo anomalies, and learn a classifier that can detect the pseudo anomalies.
Time Series Anomaly Detection:
There exists a large body of work on TSAD, for which we refer to several surveys [5, 15]. Recently, SSL has been widely studied as the main approach to address TSAD problems [32, 37]. BeatGAN [43] and RobustTAD [13] use data augmentation to enrich training data. COUTA [33] proposes three different augmentations, respectively mimicking local, contextual and collective anomalies. NeuTraL-AD [26] propose several augmentation functions that are an integral part of the learning process. AMSL [38] also uses signal transformations as data augmentation. TimeAutoAD [18] employs three different strategies to augment the training data for generating pseudo anomalies. Motivated by masked language models [11], MAD [12] investigated various masking procedures for multivariateTSAD based on a predictive pretext task. Note that only [18, 26] parameterize the augmentation procedures, but they tune those using labeled validation data.
Unsupervised Model Selection:
Unsupervised hyperparameter tuning (i.e. model selection) is nontrivial in anomaly detection due to the absence of labeled validation data [23], and the literature is slim with quite recent efforts [39, 40, 41, 42]. Specifically on SSL-based AD, a recent study [35] has revealed the impact of the choice of augmentation on SSL-based AD, leading to several works that aim to automatically search for the optimal choice [34, 36]. Their main contribution is a novel unsupervised loss, but it is not differentiable. None of the existing works addressed the model selection problem of SSL-based AD for time series data.

### v2

@@ 3 Experiments
Table 1: Detection peformance of baselines w.r.t. $F_{1}$ and AUROC on test data across six TSAD tasks (cf. Appx. C.1, Table 3). TSAP has the lowest average rank w.r.t. $F_{1}$ and AUROC, with low std. dev.
@@
OC-SVM 0.182 0.468 0.182 0.472 0.373 0.803 0.393 0.806 1.000 1.000 0.546 0.806 7.2 (3.7) 6.8 (3.3)
LOF 0.999 1.000 0.999 1.000 0.354 0.738 0.358 0.725 0.196 0.506 0.221 0.577 6.8 (3.9) 6.5 (4.4)
ARIMA 0.885 0.960 0.829 0.965 0.991 0.999 0.999 0.999 0.870 0.955 0.225 0.537 4.3 (3.2) 4.7 (3.7)
IF 0.255 0.587 0.232 0.576 0.183 0.402 0.182 0.356 0.864 0.965 0.342 0.758 8.8 (1.7) 8.7 (1.9)
MP 0.812 0.743 0.812 0.744 0.280 0.712 0.284 0.734 1.000 1.000 1.000 1.000 5.0 (3.6) 4.8 (3.4)
EncDec-LSTM 0.190 0.508 0.190 0.508 0.415 0.812 0.442 0.819 0.980 0.999 0.909 0.996 6.5 (2.0) 6.7 (1.9)
SR-CNN 0.965 0.990 0.964 0.998 0.983 0.999 0.991 0.999 0.302 0.700 0.214 0.512 5.2 (4.2) 4.8 (4.5)
USAD 0.183 0.425 0.184 0.428 0.430 0.822 0.409 0.828 1.000 1.000 1.000 1.000 5.5 (4.0) 5.7 (4.5)
NeuTraL-AD 0.211 0.732 0.263 0.679 0.561 0.868 0.526 0.862 1.000 1.000 1.000 1.000 4.2 (2.9) 3.8 (2.5)
TimeGPT 0.327 0.714 0.318 0.711 0.218 0.580 0.217 0.525 0.348 0.743 0.385 0.683 8.0 (1.9) 8.3 (1.6)
TSAP (ours) 1.000 1.000 1.000 1.000 0.973 0.999 0.991 0.998 0.889 0.969 1.000 1.000 2.3 (2.0) 2.2 (2.0)
We evaluate TSAP on six distinct TSAD tasks. Four of these are conducted in a controlled environment, using the 2017 ECG PhysioNet Challenge dataset [9]. The remaining two are natural anomalies, derived from the CMU Motion Capture (MoCap) dataset. In the former, the anomaly types are manually injected in $\mathcal{D}_{\mathrm{test}}$ based on the types discussed in Sec. 2.1. For the latter, the anomaly types in $\mathcal{D}_{\mathrm{test}}$ are unknown and it is the goal of TSAP to find the type that yields best alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ (part of $\mathcal{D}_{\mathrm{test}}$), expressed by $\mathcal{L}_{\mathrm{val}}$. See Appx. C for further details.
Quantitative Results
@@
Table 1 shows the detection results for these six TSAD tasks. TSAP ranks the best overall in both $F_{1}$ and AUROC, showing that detector $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$, trained through the alternating mechanism of TSAP, is able to generalize to unseen, unlabeled anomalies in $\mathcal{D}_{\mathrm{test}}$. While competing methods perform well on certain tasks, they lack consistency across all TSAD tasks. LOF, for instance, thrives on PhysioNet A and B due to its abrupt, manually injected anomalies, but fails on MoCap, where anomalies emerge more gradually as activities transition naturally. SR-CNN and ARIMA show similar behavior. See Appx. B.2B.3 for model/baseline configuration details.
Reconstruction-based methods like EncDec-LSTM and USAD struggle on PhysioNet due to high variability among inliers, yet excel on MoCap, which exhibits consistent, near-periodic patterns. NeuTraL-AD, despite its augmentation-based approach, struggles with PhysioNet’s real-world ECG signal variability, suggesting its augmentation functions lack robustness in noisy conditions. We remark that only TSAP provides robust and consistent performance across all TSAD tasks.
Qualitative Results
@@
Figure 2: Tuning continuous augmentation hyperparameter(s) with TSAP. Top: For Platform anomalies at true level (red dashed line), various initializations converge near the true value (left) by minimizing val. loss (center), which leads to high detection performance (right). Bottom: Multiple continuous hyperparameters (level and length) converge near true values (left) through minimizing val. loss (center), resulting in high AUROC (right). Diverged results correspond to high val. loss, aiding in rejecting low-performance models.
@@
Figure 3: Tuning the discrete hyperparameter (anomaly type) with TSAP. Top: For true Trend anomaly (black), val. loss favors Trend (green) and Mean shift (blue), both achieving high AUROC (middle) and strong resemblance (right), while rejecting Platform (orange) with poor performance. Bottom: For Jump anomalies in MoCap A with unknown type (black), val. loss favors Platform (purple), leading to high AUROC (middle) and similarity to the true anomaly (right), while rejecting Frequency (red) with poor performance.
Key to TSAP’s consistent performance is $\mathcal{L}_{\mathrm{val}}$ through which TSAP automatically learns the augmentation hyperparameters $\mathbf{a}$. Once $\mathbf{a}$ is determined, the task reduces to a supervised learning problem. Next, we show that TSAP not only effectively tunes the continuous augmentation hyperparameters $\mathbf{a}$, but also that $\mathcal{L}_{\mathrm{val}}$ guides the accurate selection of the discrete hyperparameter (i.e. anomaly type). Consider PhysioNet A, where we aim to tune the continuous hyperparameter $a$, i.e. level, of the Platform anomalies present in $\mathcal{D}_{\mathrm{test}}$. That is, the level in $\mathcal{D}_{\mathrm{test}}$ is fixed and tuning aims to estimate its value using TSAP while the other hyperparameters (location, length) are randomized. Fig. 3 (top) shows TSAP’s estimation process for different initializations of $a$. We observe that the initialization for $a\in\{-0.4,0.6\}$ leads to the true $a$$=$$0.2$ (left). Simultaneously, $\mathcal{L}_{\mathrm{val}}$ drops substantially once TSAP has arrived at the true $a$ (center). This is also reflected in the performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ which soars upon estimation of the true $a$ (right). Conversely, the initialization for $a\in\{0.8,-0.8\}$ leads to a high $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$. Indeed, the performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ now suffers from poor alignment. For PhysioNet B, we estimate both level and length while location is randomized. Fig. 3 (bottom) demonstrates TSAP’s ability to accurately estimate the level and length. Further, Fig. 3 (top) showcases TSAP’s ability to perform discrete hyperparameter selection. Here, TSAP has been initialized and trained with three anomaly types (Mean shift, Platform, Trend) on PhysioNet C (true anomaly type: Trend). $\mathcal{L}_{\mathrm{val}}$ indicates a misalignment between the Platform and Trend types (left), also reflected in the AUROC of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ (center). Note how the Mean shift anomaly type has a low $\mathcal{L}_{\mathrm{val}}$ at the end of the training epochs, reflected in the high AUROC on $\mathcal{D}_{\mathrm{val}}$. This shows that the true underlying anomaly type is not necessarily the only type that yields high alignment, and in turn a high-performing detector. In MoCap datasets, where anomaly types are a priori unknown, TSAP is initialized with different augmentation types (Frequency and Platform) to perform discrete hyperparameter selection. Fig. 3 (bottom) highlights its effectiveness as $\mathcal{L}_{\mathrm{val}}$ clearly prefers one type over the other. Indeed, the natural anomalies defined by jumping signals in MoCap A have close resemblance to platform anomalies. See Appx. D for additional results and ablation studies.